OpenStudy (evonhowell):

When solving systems of equations you should find the easiest way to use based on the equation that you are working with. For this equation the easiest I could find was elimination 1st we need to salve for “x” 6x – 4y = 54 –9x + 2y = –69 You need to multiply by 2 on this equation and then combine the two equations. 3x = -15 X=5 2nd we need to solve for “y” 6x - 4y = 54 -9x + 2y = -69 6y= 5 Y= 1.2 Elimination IS the easiest to use.

OpenStudy (unklerhaukus):

these are your simultaneous equations? \[6x – 4y = 54\qquad(i)\] \[–9x + 2y = –69\qquad(ii)\]

OpenStudy (anonymous):

\[6x-4y=54\]\[-9x+2y=-69\]\[6x-4y+2(-9x+2y)=54+2(-69)\]\[6x-4y-18x+4y=54-138\]\[-12x=-84\]\[x=\frac{ -84 }{ -12 }=7\]

OpenStudy (evonhowell):

all it says is math processing error

OpenStudy (unklerhaukus):

reload

OpenStudy (evonhowell):

ok I did

OpenStudy (evonhowell):

yes Un kleRhakkus

OpenStudy (evonhowell):

in reply to your first reply

OpenStudy (unklerhaukus):

Aylin has solved using the method you suggested but as you can see you made some mistakes in your working

OpenStudy (evonhowell):

Yes I did thank you guys very much :)

OpenStudy (unklerhaukus):

i might suggest a slight variation \[6x–4y=54\qquad(i)\]\[–9x+2y=–69\qquad(ii)\] divide (i) by 2 \[3x–2y=27\qquad(iii)\] now add (ii) and (iii) \[(–9x+2y)+(3x–2y)=(–69)+(27)\] \[–6x=-42\]\[x=\frac{-42}{-6}=7\]

OpenStudy (anonymous):

Dividing by 2 is probably easier. I multiplied the second instead so it would be easier for EvonHowell to follow what I did and see where his mistake was made.

OpenStudy (evonhowell):

well, Thank you guys soooo much :)

OpenStudy (anonymous):

Glad to help. :)

OpenStudy (evonhowell):

:)